# Robotic Ultrasound-Guided Femoral Artery Reconstruction of Anatomically-Representative Phantoms

### v1

@@ III-B2 Vascular Segmentation NetworkVideos
We propose a deep learning network to segment vessels in US videos by extending a traditional single-frame segmentation network into a video segmentation pipeline. The network processes preceding frames using encoders to extract key and value features, which are stored in a memory module. This memory module retains the value features alongside their corresponding key features, which are subsequently accessed and integrated with query feature maps through an attention-based module. The fused features generated by this process are then passed to a decoder, which produces the final segmentation mask. An overview of the network architecture is presented in Fig. 5. The key components of the network are explained below:
@@
Memory: The extracted key and value features are stored in a memory. During the segmentation of the query frame, temporal information must be retrieved from memory. This retrieval is accomplished by matching the key feature maps from the memory and the query frames, denoted as $\boldsymbol{{K}}_{M}^{s_{i}}$ and $\boldsymbol{{K}}_{Q}$, respectively. To optimize this process, we leverage temporal prior information as in [35], which assumes that the content in the query frame is more likely to appear in similar pixel locations within the preceding (memory) frames. This operation is implimented as correlation operation:
@@
The probability maps represent the attention weights, which are used to weigh both the spatial and temporal information stored in the memory. The final output tensor from the memory can be expressed as
$\begin{split}\tilde{\boldsymbol{{V}}}_{M}(x,y)=\sum_{s_{i}\in\mathcal{S}}\sum_{(n,m)\in\mathcal{R}}\boldsymbol{\mathcal{P}}_{nm}^{s_{i}}(x,y)\boldsymbol{{V}}_{M}^{s_{i}}(x+n,y+m),\end{split}$ (3)
which represents the weighted sum of all memory values $\boldsymbol{V}_{M}^{s_{i}}$ across both the temporal domain and spatial domain.
Fusion: The fusion module combines the $\boldsymbol{V}_{Q}$ and $\tilde{\boldsymbol{{V}}}_{M}$ to construct a relevant input for the decoder to carry out the segmentation of the query frame. To achieve this, an attention branch is assigned to each fusion input, as in [35]. These attention branches, equipped with sigmoid-activated gates, regulate the flow of information by controlling which features - either from the query or from memory - are passed to the decoder. The fusion module can be mathematically expressed as
@@
Dataset: US images were acquired using the Clarius C3 HD scanner, with a hydrogel block placed over only the old CIRS phantom. This block matches the CIRS phantom in length and width but a varying depth. This is done to keep the physics of ultrasound image same as representative dataset from developed phantoms. Scans were performed along all four vessels at a 100mm depth, capturing a total of 4,802 images across 10 continuous scan sequences. Vessel structures were manually segmented by two annotators using the VIA tool [36]. The dataset was partitioned into training, validation, and testing sets in an 80:10:10 ratio (3,838, 476, and 488 images, respectively), ensuring proportional sequence representation across all subsets. Additional dataset statistics are provided in Table I.
@@ III-B3 Vascular Centerline Estimation
Following robotic US image acquisition and vessel segmentation, postprocessing is implemented to mitigate noise and artifacts. False positive detections are eliminated by excluding centroids located beyond ±85 pixels from the image center, corresponding to a 45mm width threshold (based on a 0.2694 mm/pixel scale factor at 100mm scanning depth). This threshold is empirically derived through analysis of femoral artery characteristics across the five patient models, where the maximum observed femoral artery bifurcation width was 39.68mm bilaterally. To account for variability, a 12.5% margin is incorporated, adjusting the threshold from 40 to 45mm. Contour detection is applied to the refined segmentation mask, retaining only nonzero-area contours. An ellipse is fitted to each valid contour to approximate vessel geometry, recording its centroid and area. Centroids are then transformed from pixel to 3D world coordinates using the US calibration matrix, with their stacked positions forming the reconstructed artery centerline. CT scans of both CIRS phantoms are acquired, and two distinct CAD models of the bifurcated vessels are generated. The vessels’ centerline is then extracted using 3D Slicer and the Vascular Modeling Toolkit [37], discretized at 1mm intervals to serve as the ground truth.
@@ III-B4 Evaluation Metrics
To evaluate reconstruction accuracy, the detected centerline is aligned to the corresponding ground truth using ICP. Spatial accuracy is quantified by computing the Hausdorff distance and $L_{2}$ norm between the reconstructed centroids and the ground truth centerline. The Hausdorff distance $H(A,B)$ between two point sets $A$ and $B$ is defined as:
$H(A,B)=\max\{\sup_{a\in A}\inf_{b\in B}||a-b||,\sup_{a\in A}\inf_{b\in B}||b-a||\},$ (5)
where |||||$\cdot$||| | ⋅ | | is the Euclidean distance. The Hausdorff distance in this case is computed from the reconstructed phantom centroids to the discretized ground truth centerline, since the reconstructed centerline is more sparse compared to the ground truth. The $L_{2}$ norm is computed as:
$L_{2}=\sqrt{\sum^{N}_{i=1}||p_{i}-q_{i}||^{2}},$ (6)
@@
Results are reported for both postprocessed detected centroids, and manually filtered centroids that exclude points not corresponding to the vessel. This evaluation does not assess the quality of US image segmentation, which is evaluated separately, but rather focuses on quantifying potential distortions in the 3D reconstruction of the artery centerline using our approach
@@ IV-A Vascular Segmentation in US Images
The network was trained on an NVIDIA GeForce RTX 4090 GPU (24GB) using Python $3.8$, PyTorch $2.4$, and CUDA $12.1$. The Adam optimizer was initialized with a learning rate of $2\times 10^{-4}$, adjusted by a factor of $0.8$ upon plateau detection, with a minimum threshold of $1\times 10^{-9}$. A batch size of $16$ was used. The input consisted of video sequences incorporating the previous $3$ frames ($s_{f}=3$). The TransUNet model was initially trained from scratch, with pretrained encoder weights subsequently transferred to the video-based model. A correlation window size of $5$ and a memory sequence length of $4$ were used for feature fusion. US images and corresponding ground truth labels were resized to $256\times 256$ pixels using bilinear and nearest-neighbor interpolation, respectively. Data augmentation techniques included random horizontal flipping, contrast adjustment via color jitter, rotations within ±$\pm 30^{\circ}$, and gamma correction with values randomly sampled between $1.0$ and $2.5$. The network was optimized using a joint Dice and Binary Cross-Entropy loss function for a maximum of $50$ epochs.
The network’s segmentation performance was assessed using the Dice Coefficient (DC) and Intersection over Union (IoU). On the training set, the model achieved a DC of 91.28% and an IoU of 83.95%, while on the test set, it attained a DC of 89.21% and an IoU of 80.54%, demonstrating good segmentation accuracy and generalization. The minimal performance disparity between training and test sets suggests that the model effectively mitigates overfitting. Figure 6(a) presents qualitative results from the test set, illustrating predicted segmentation masks overlaid on input images. The close correspondence between predicted and ground-truth masks further validates the model’s robustness in arterial segmentation. Due to the absence of standardized benchmark datasets for vascular segmentation, direct comparative analysis with existing methodologies presents inherent challenges. Nevertheless, for contextual reference, Morales et al. [38] reported an IoU of 77.9% on a medical phantom using their proposed technique, and an average IoU of 63.8% on pig US data.
@@ IV-B Vascular Centerline Reconstruction
Examples of the reconstructed centerline overlaid on the ground truth centerline for the bifurcated arteries of the first (older) phantom are shown in Figure 6(b)-(c). The figure includes results both before and after manually removing outliers corresponding to erroneously detected center points by the network. Additionally, violin plots illustrating the $L_{2}$ errors are shown in Figure 7, with corresponding metrics such as the mean, standard deviation, and Hausdorff distance reported in Table II.
@@
TABLE II: Table presenting the mean error, standard deviation, and Hausdorff metrics for the different CIRS phantoms (PH1 and PH2) and patients (P1 through P5) with and without removal of erroneous datapoints, and the corresponding average values per CIRS phantom.
Mean $L_{2}$ Err. (mm) Std (mm) Hausdorff (mm)
PH 1 P1 Unfiltered 2.957 6.410 41.77
Filtered 0.741 0.463 3.046
P2 Unfiltered 3.503 7.613 58.67
Filtered 1.105 0.788 6.059
P3 Unfiltered 3.819 7.026 31.96
Filtered 0.890 0.815 5.241
P4 Unfiltered 1.711 4.777 49.56
Filtered 0.955 0.821 3.820
P5 Unfiltered 1.223 2.449 21.71
Filtered 0.887 0.629 3.647
Avg Unfiltered 2.642 5.655 40.73
Filtered 0.915 0.703 4.362
PH 2 P1 Unfiltered 2.471 4.950 15.10
Filtered 0.746 0.836 6.216
P2 Unfiltered 4.933 7.923 19.28
Filtered 1.629 1.038 6.355
P3 Unfiltered 2.358 4.349 37.68
Filtered 1.343 1.183 6.922
P4 Unfiltered 6.587 6.297 34.71
Filtered 1.645 1.035 4.743
P5 Unfiltered 1.533 2.907 22.72
Filtered 1.151 1.249 5.499
Avg Unfiltered 3.576 5.285 25.89
Filtered 1.302 1.068 5.947
@@
The 3D points reconstructed from the segmented axial US images show strong spatial alignment with the ground truth centerlines extracted from CT scans. Similar consistency was observed in results obtained from the second CIRS model, demonstrating the robustness of our proposed method. These findings indicate that the transformation from the segmented US space to the 3D world frame preserves spatial accuracy without introducing distortions - a critical requirement for reconstructing small-scale vascular structures such as the femoral artery, which can have diameters as small as 4 mm.
@@
The importance of filtering out erroneous centroids is best illustrated from Figure 6. The unfiltered data shows a naturally wider spread compared to that of the filtered data, making it challenging to adequately assess the reconstruction quality. The Hausdorff errors are understandably larger than the $L_{2}$ errors, as the Hausdorff metric in this context reports the largest $L_{2}$ error within the dataset (as reflected by the maximum values in the violin plots). The average filtered $L_{2}$ and Hausdorff errors for the first vascular phantom (0.915 and 4.362mm, respectively) are smaller than those of the second phantom (1.302 and 5.947mm), as shown in Table II. This discrepancy can be attributed to the poorer quality of the US images for the second phantom, leading to false positives in the segmentation. It is important to highlight that the network was trained exclusively on the first CIRS model. While both phantoms are theoretically identical, their appearance under US differs in practice due to material variations and imaging artifacts. Additionally, the smaller bifurcated vessel in the second phantom, with its reduced diameter, was frequently missed by the network for both phantoms, as it became more difficult to visualize it, especially with phantoms of patients with higher BMI. This resulted in a more sparse detection of centroids for that vessel. The issue may have been further exacerbated by the use of a convex US probe. While linear probes are typically preferred for vascular scanning, a convex probe was used due to availability constraints. As a result, probe contact with the phantom did not always ensure full transducer-surface coupling, leading to suboptimal imaging. This effect was particularly pronounced in scenarios where the natural body contours would necessitate different probe orientations, even for a skilled clinician.
@@ V Conclusions and Future Work
While robotic US for vascular access has seen significant advancements, prior studies have not validated vessel 3D reconstruction on anatomically representative phantoms. In this work, we present a robotic US-guided system for femoral artery reconstruction on a diverse set of 5 patient-specific phantoms. Our approach achieves high-fidelity vessel reconstructions even in complex phantoms with irregular topology, including deeper arteries in patients with higher BMI. Additionally, we adapted a video-based segmentation network for axial artery segmentation in US video sequences. The reconstructed artery centerline exhibited strong agreement with patient-specific CT data, the clinical gold standard for vascular imaging. These results represent a step forward in enhancing robotic US-assisted vascular access procedures, improving their accuracy and potential for clinical translation.

### v2

@@ III-B2 Vascular Segmentation Network
We propose a deep learning network to segment vessels in US videos by extending a traditional single-frame segmentation network into a video segmentation pipeline. The network processes preceding frames using encoders to extract key and value features, which are stored in a memory module. Since US images often exhibit speckle noise, low contrast, and transient artifacts, the memory module leverages temporal context from preceding frames to help maintain spatial continuity in the segmentation. The memory module retains the value features alongside their corresponding key features, which are subsequently accessed and integrated with query feature maps through an attention-based module. The fused features generated by this process are then passed to a decoder, which produces the final segmentation mask. An overview of the network architecture is presented in Fig. 5. Key components of the network are explained below:
@@
Memory: The extracted key and value features are stored in a memory module. During the segmentation of the query frame, temporal information must be retrieved from memory. This retrieval is accomplished by matching the key feature maps from the memory and the query frames, denoted as $\boldsymbol{{K}}_{M}^{s_{i}}$ and $\boldsymbol{{K}}_{Q}$, respectively. To optimize this process, we leverage temporal prior information as in [35], which assumes that the content in the query frame is more likely to appear in similar pixel locations in the preceding (memory) frames. This operation is implemented as a correlation operation:
@@
The probability maps represent attention weights, which are used to weigh both spatial and temporal information stored in the memory. The final output tensor from the memory can be expressed as
$\tilde{\boldsymbol{{V}}}_{M}(x,y)$ $=\sum_{s_{i}\in\mathcal{S}}\sum_{(n,m)\in\mathcal{R}}\boldsymbol{\mathcal{P}}_{nm}^{s_{i}}(x,y)$ (3)
$\quad\times\boldsymbol{{V}}_{M}^{s_{i}}(x+n,y+m),$
which represents the weighted sum of all memory values $\boldsymbol{V}_{M}^{s_{i}}$ across both temporal domain and spatial domains.
Fusion: The fusion module combines $\boldsymbol{V}_{Q}$ and $\tilde{\boldsymbol{{V}}}_{M}$ to form the decoder input needed for segmenting the query frame. To achieve this, an attention branch is assigned to each fusion input, as in [35]. These attention branches, equipped with sigmoid-activated gates, regulate the flow of information by controlling which features - either from the query or from memory - are passed to the decoder. The fusion module can be mathematically expressed as
@@
Dataset: US images were acquired using the Clarius C3 HD scanner, with a hydrogel block placed over only the old CIRS phantom. The block matched the phantom’s length and width but had a variable thickness, ensuring that the acoustic propagation characteristics were consistent with those of the patient-specific phantoms. Scans were performed along all four vessels at a 100mm depth, capturing a total of 4,802 images across 10 continuous scan sequences. Vessel structures were manually segmented by two annotators using the VIA tool [36]. The dataset was partitioned into training, validation, and testing in an 80:10:10 ratio (3,838, 476, and 488 images, respectively), ensuring proportional sequence representation across all subsets. Additional dataset information is provided in Table I.
@@ III-B3 Vascular Centerline Estimation
Following robotic US image acquisition and vessel segmentation, post-processing is applied to suppress noise and reduce artifacts. False positive detections are eliminated by excluding centroids located beyond ±85 pixels from the image center, corresponding to a 45mm width threshold (based on a 0.2694 mm/pixel scale factor at 100mm scanning depth). This threshold is empirically derived through analysis of femoral artery characteristics across the five patient models, where the maximum observed femoral artery bifurcation width was 39.68mm bilaterally. To account for variability, a 12.5% margin is incorporated, adjusting the threshold from 40 to 45mm. Contour detection is applied to the refined segmentation mask, retaining only contours with non-zero area. An ellipse is fitted to each valid contour to approximate vessel geometry, recording its centroid and area. Centroids are then transformed from pixel to 3D world coordinates using the US calibration matrix, with their stacked positions forming the reconstructed artery centerline. CT scans of both CIRS phantoms are acquired, and two distinct CAD models of the bifurcated vessels are generated. The vessels’ centerline is then extracted using 3D Slicer and the Vascular Modeling Toolkit [37], discretized at 1mm intervals to serve as the ground truth.
@@ III-B4 Evaluation Metrics
To evaluate the reconstruction accuracy, the detected centerline is aligned to the corresponding ground truth using ICP. Spatial accuracy is quantified by computing the Hausdorff distance and $L_{2}$ norm between the reconstructed centroids and ground truth centerline. The Hausdorff distance $H(A,B)$ between two point sets $A$ and $B$ is defined as:
[…]
where $||\cdot||$ is the Euclidean distance. The Hausdorff distance in this case is computed from the reconstructed phantom centroids to the discretized ground truth centerline, since the reconstructed centerline is more sparse compared to the ground truth. The $L_{2}$ norm is computed as:
[…]
@@
Results are reported for both postprocessed detected centroids, and manually filtered centroids that exclude points not corresponding to the vessel. This evaluation does not assess the quality of US image segmentation per se, which is evaluated separately, but rather focuses on quantifying potential distortions in the 3D reconstruction of the artery centerline using our approach.
@@ IV-A Vascular Segmentation in US Images
The network was trained on an NVIDIA GeForce RTX 4090 GPU (24GB) using Python 3.8, PyTorch 2.4, and CUDA 12.1. The Adam optimizer was initialized with a learning rate of $2\times 10^{-4}$, adjusted by a factor of 0.8 upon plateau detection, with a minimum threshold of $1\times 10^{-9}$. A batch size of 16 was used. The input consisted of video sequences incorporating the previous 3 frames ($s_{f}=$ 3). The TransUNet model was initially trained from scratch, with pretrained encoder weights subsequently transferred to the video-based model. A correlation window size of 5 and a memory sequence length of 4 were used for feature fusion. US images and corresponding ground truth labels were resized to $256\times 256$ pixels using bilinear and nearest-neighbor interpolation, respectively. Data augmentation techniques included random horizontal flipping, contrast adjustment via color jitter, rotations within $\pm 30^{\circ}$, and gamma correction with values randomly sampled between 1.0 and 2.5. The network was optimized using a joint Dice and Binary Cross-Entropy loss function for a maximum of 50 epochs.
The network’s segmentation performance was assessed using the Dice Coefficient (DC) and Intersection over Union (IoU). On the training set, the model achieved a DC of 91.28% and an IoU of 83.95%, while on the test set, it attained a DC of 89.21% and an IoU of 80.54%, demonstrating good segmentation accuracy and generalization. The minimal performance disparity between training and test sets suggests that the model effectively mitigates overfitting. Figure 6(a) presents qualitative results from the test set, illustrating predicted segmentation masks overlaid on input images. The close correspondence between predicted and ground-truth masks further validates the model’s robustness in arterial segmentation. Due to the absence of standardized benchmark datasets for vascular segmentation, direct comparative analysis with existing methodologies presents inherent challenges. Nevertheless, for contextual reference, Morales et al. [38] reported an IoU of 77.9% on a medical phantom using their proposed technique, and an average IoU of 63.8% on pig US data. The inference time on the same GPU used for training is 6ms per 3-frame video sequence.
@@ IV-B Vascular Centerline Reconstruction
The robotic scanning protocol was completed in 1 min and 23 seconds for each of the 10 experimental cases. Examples of the reconstructed centerline overlaid on the ground truth centerline for the bifurcated arteries of the first (older) phantom are shown in Figure 6(b)-(c). The figure includes results both before and after manually removing outliers corresponding to erroneously detected center points by the network. Additionally, violin plots illustrating the $L_{2}$ errors are shown in Figure 7, with corresponding metrics such as the mean, standard deviation, and Hausdorff distance reported in Table II. Values in red denote the highest metrics, where values in blue denote the lowest. The number of detected centroids with and without manual filtering is also reported in Table II.
@@
TABLE II: Mean error, standard deviation, Hausdorff, and number of detected centroid points for all CIRS phantoms (PH1 and PH2) and patients (P1 through P5) with and without removal of outliers. All values (except for points #) are in $m$. Values in red denote highest metrics, values in blue denote the lowest.
Mean $L_{2}$ Err. Std Hausdorff Points #
PH 1 P1 Unfiltered 2.957 6.410 41.77 2659
Filtered 0.741 0.463 3.046 2340
P2 Unfiltered 3.503 7.613 58.67 2591
Filtered 1.105 0.788 6.059 2305
P3 Unfiltered 3.819 7.026 31.96 2110
Filtered 0.890 0.815 5.241 1719
P4 Unfiltered 1.711 4.777 49.56 2632
Filtered 0.955 0.821 3.820 2482
P5 Unfiltered 1.223 2.449 21.71 3339
Filtered 0.887 0.629 3.647 3271
Avg Unfiltered 2.642 5.655 40.73 2666
Filtered 0.915 0.703 4.362 2423
PH 2 P1 Unfiltered 2.471 4.950 15.10 2045
Filtered 0.746 0.836 6.216 1770
P2 Unfiltered 4.933 7.923 19.28 2017
Filtered 1.629 1.038 6.355 1678
P3 Unfiltered 2.358 4.349 37.68 2491
Filtered 1.343 1.183 6.922 2324
P4 Unfiltered 6.587 6.297 34.71 1168
Filtered 1.645 1.035 4.743 957
P5 Unfiltered 1.533 2.907 22.72 2724
Filtered 1.151 1.249 5.499 2667
Avg Unfiltered 3.576 5.285 25.89 2089
Filtered 1.302 1.068 5.947 1879
@@
The 3D points reconstructed from the segmented axial US images show strong spatial alignment with the ground truth centerlines extracted from CT scans. Similar consistency was observed in results obtained from the second CIRS model, demonstrating the robustness of our proposed method. These findings demonstrate that the transformation from the segmented US space to the 3D world frame preserves spatial accuracy without introducing distortions - a critical requirement for reconstructing small-scale vascular structures, such as the femoral artery, with diameters as small as 4mm.
@@
The importance of filtering out erroneous centroids is best illustrated in Figure 6(b)-(c). The unfiltered data exhibits a naturally wider spread of points than the filtered data, making it difficult to reliably assess reconstruction quality in isolation. The Hausdorff errors are understandably larger than the $L_{2}$ errors, as the Hausdorff metric in this context reports the largest $L_{2}$ error within the dataset (as reflected by the maximum values in the violin plots). The average filtered $L_{2}$ and Hausdorff errors for the first vascular phantom (0.915 and 4.362mm, respectively) are smaller than those for the second phantom (1.302 and 5.947mm), as shown in Table II. This discrepancy likely arises from the lower quality of the US images for the second phantom, which increased the occurrence of false positives in the segmentation. It is important to highlight that the network was trained exclusively on the first CIRS model. While both phantoms are theoretically identical, their appearance under US differs in practice due to material variations and imaging artifacts. Additionally, the smaller bifurcated vessel in the second phantom, with its reduced diameter, was frequently missed by the network due to limited visibility - a challenge further exacerbated in higher-BMI patient phantoms - resulting in sparser centroid detections for that vessel. The use of a convex US probe may have further contributed to this issue. Probe contact with the phantom did not always ensure full transducer-surface coupling, leading to suboptimal imaging. This effect was particularly pronounced in scenarios where the natural body contours would necessitate different probe orientations, even for a skilled clinician. A linear probe, commonly used for vascular access procedures, could improve acoustic coupling and, in turn, enhance segmentation performance.
@@ V Conclusions and Future Work
While robotic US for vascular access has seen significant advancements, prior studies have not validated vessel 3D reconstruction on anatomically representative phantoms. In this work, we present a robotic US-guided system for femoral artery reconstruction on a diverse set of 5 patient-specific phantoms. Our approach achieves high-fidelity vessel reconstructions even in complex phantoms with irregular topology, including deeper arteries in patients with higher BMI. Additionally, we adapted a video-based segmentation network for axial artery segmentation in US video sequences. The reconstructed artery centerline exhibited strong agreement with patient-specific CT data, the clinical gold standard for vascular imaging. These results represent a step forward in enhancing robotic US-assisted vascular access procedures, strengthening their potential for clinical translation.